### v1

# Lyapunov Dynamics in Entangled Biphoton Spectroscopy

Abstract
Recent advances in quantum light spectroscopy highlight the potential of using entangled photons as a sensitive probe for many-body dynamics and material correlations. However, a comprehensive theory to explain experimental results remains elusive, primarily due to the complexity of the Hilbert space and the intricate interactions and nonlinearities inherent in material systems. In this work, we introduce a tractable model based on a finite-sized correlation matrix governed by a bilinear bosonic Hamiltonian, enabling efficient simulations through Gaussian-preserving dynamics. We apply this framework to compute the output joint spectral intensity (JSI) and von Neumann entropy of frequency-entangled biphotons, and find close agreement with experimental observations in empty microcavities. Our results reveal the emergence of off-diagonal spectral correlations that can be interpreted as irreversible decay of cavity excitations into the biphoton continua. This approach offers a powerful theoretical tool for interpreting quantum spectroscopic data and paves the way for probing more complex light-matter interactions in materials.
I Introduction
A growing literature in theoretical and experimental quantum light spectroscopy techniques reveals that quantum light can be used as a sensitive tool to probe the electronic and vibrational properties of materials [1, 2, 3, 4, 5, 6, 7, 8, 9]. The ability to perform spectroscopic measurements using low-photon fluxes becomes significant when exploring the energy conversion dynamics in photovoltaic materials, microscopic phenomena in biological systems, and any other such systems where understanding the many-body correlations and fluctuations reveal richer information than classical spectroscopic techniques. Traditionally, quantum spectroscopic experiments look at the changes in entanglement of photons as they interact with the many degrees of freedom of the material considered. Recent progress in this direction employs frequency-entangled biphotons as the probe as seen in Malatesta et al. [10] and Moretti et al. [11] These experiments use a process called Spontaneous Parametric Down Conversion (SPDC) where a pump photon, incident on a nonlinear crystal, generates two channels of photons, conventionally called the signal and idler channels. This process conserves the frequency and momentum of the pump and produces the two channels with a distributed amplitude along a frequency range. We refer to this amplitude as the Joint Spectral Amplitude (JSA). However, in photon-counting experiments, one usually has to deal with a delay between two consecutive coincidence detection events. As a result, the data obtained is effectively a time-integrated probability amplitude over the delay period. This is known as the Joint Spectral Intensity (JSI), which is usually the observable in quantum spectroscopic experiments.
In order to realize a viable platform to conduct these experiments, one needs to fabricate strong light-matter interactions. This is achieved by placing materials within optical microcavities. These not only have the ability to amplify electric fields, but also provide quantized electromagnetic modes (EM) that interact with materials. These quantized modes enable us to probe the dynamics between multiple excitons and provide a comprehensive understanding of the many-body correlations that exist in materials. These EM modes interacting with material excitations form strongly-interacting light-matter hybrid particles called polaritons that offer a testbed to realize a plethora of exotic physics. These include Bose-Einstein Condensation of polaritons [12, 13, 14], superfluidity and vortex formations [15, 16], nonlinear effects such as optical bistability [17, 18], four-wave mixing [19], and soliton formation [20, 21]. Moreover, in specially engineered photonic lattices, polaritons can exhibit topological protection, similar to the quantum Hall effect or topological insulator phases [22, 23]. In two-dimensional planar materials such as and graphite, phonon polaritons can exhibit hyperbolic dispersion [24, 25, 26], enabling deep subwavelength confinement. Since these polaritons represent strong light-matter interactions, they form the perfect spectroscopic tool to probe many-body correlations using single- and biphoton sources.
Interpreting experimental results requires a robust theoretical framework in quantum spectroscopy. Although ample theoretical development is available in the literature, they fall short in reproducing experimental observations and accurately explaining the physics behind polariton-induced scattering. Bittner et al. [27] showed how radiative cascade effects and biphoton scattering among molecular dimers lead to a generation in the output entanglement entropy of biphoton states. In this, they treat the dimer as coupled excitons with anharmonic interactions embedded within a continuum of photons and show, using diagrammatic expansions, that the coupling strength and repulsions can be correlated to the output entropy. In addition, Li et al. [28, 29] also laid a similar framework to establish photon-photon correlations using the Dicke model and the input-output theory [30, 31] in a Hong-Ou-Mandel apparatus [32]. In this apparatus, only one of the photons of the biphotons produced from the SPDC process interacts with a resonant medium. As this photon passes through the medium, a transmission function is introduced that is multiplied by the input to give the output photon state. All many-body dynamics of the resonant medium is then traced back by studying the pole structure of this function and analyzing the dynamics of the resulting polariton branches. While these developments provide detailed frameworks for describing many interesting and non-linear photophysics, their numerical simulations become quite challenging owing to an exponentially growing Hilbert space from the many biphoton modes. For example, for a biphoton state interacting with a cavity-polariton under the Jaynes-Cummings Hamiltonian, the Hilbert space is of the order of where is the number of signal/idler photon modes. Moreover, it becomes imperative to describe the mapping between the input and output modes of the system when attempting to reconcile with tradional experimental practices. In particular, the development of a numerically-tractable framework that directly connects the input and output JSIs can prove to be a fruitful foundation upon which more complicated photophysics can be built.
To address these gaps, we present a simple model based on the dynamics of correlation matrices. We consider the governing Hamiltonian to comprise bilinear bosonic operators. Consequently, the time-evolution unitary operator, , becomes a gaussian preserving map. This allows us to translate our dynamics into the first moments and covariance matrices of the problem. As a result, we can perform numerical simulations using finite matrices and arrive at convergent solutions to the output JSI. This can be done by solving the resultant Lyapunov equations of the covariance matrix and extract the output correlations between the signal and idler photons. These correlations represent the output JSA which can then be compared to the input.
Additionally, we use Møller operators to connect the input, interacting, and output variables of our experiment, eliminating the need for discrete time gating and yielding a framework that is commensurate with experimental practice. These operators provide an elegant method to connect the input and output covariance matrices that are mediated by the Lyapunov equations. Other observables, like the purity of the output biphoton state, can also be seamlessly extracted from the Wigner function expressed in terms of the covariance matrix. This framework also allows us to easily extend our study to the Tavis-Cummings limit, enabling a deeper probe of material systems. We demonstrate that with these considerations, our model predicts an output JSI that bears close resemblance with experimental observations of biphotons interacting with an empty microcavity.
II Theoretical Framework
II.1 Hamiltonian
Consider a coupled-oscillator model of frequency-entangled bath photons interacting with cavity modes and material excitations according to the following Hamiltonian,
where are the cavity modes, are the biphoton signal/idler operators, and the are operators of the material modes that are assumed to be bosonic. The first 3 terms of Eq. (II.1) are the free Hamiltonians of the cavity, signal/idler photons and the material respectively. The fourth term describes the bilinear coupling between the signal/idler modes with the photons and the last term describes the coupling between the cavity and material modes. The coupling terms obey the first Markov and rotating wave approximations. The former assumes a uniform coupling strength and the latter removes rapid counter-rotating terms.
We assume that the signal and idler photons begin in an entangled state given by,
where is the joint spectral amplitude of the input photons. This two-dimensional correlation function quantifies the degree of entanglement between the signal and idler photon states.
Fig. 1 shows an example of an experimental input and output biphoton JSI for empty microcavities. The input JSA represents a spectrally entangled state that is generated in a Type-I -Barium Borate (BBO) crystal phase-matched for SPDC close to the degeneracy at the pump wavelength of nm [10]. This dictates the signal/idler dispersion in Eq. (II.1). Several different input states have also been experimentally produced as demonstrated in Moretti et al. [11] Similarly, the output state can be defined as,
The subscripts ”in” and ”out” in the operators act on the input and output vacuum respectively. According to the input-output formalism [30, 31], the output and the input operators are related to the coupled cavity mode as,
This form of the input-output relation assumes the transmission of the input photons completely onto the output without reflections back to the input vacuum. Since these modes admit asymptotic solutions in time as the initial and final states, we can relate the freely evolving input and output modes from Eqs. (2), (3) with the interacting modes in Eq. (II.1) using Møller operators [33].
where and are the free and interacting Hamiltonians, and are the free and interacting modes respectively. More generally, we can define a vector of modes as,
and define the Møller operators for these modes as
where the specified Hamiltonians are marked in Eq. (II.1).
II.2 Governing Equation of the Model
The Hamiltonian in Eq. (II.1) can be numerically intractable with current computational capabilities. In order to make this problem solvable, we note that all the operators in the Hamiltonian are bilinear. As a result, the unitary operator is a gaussian-preserving map, allowing us to reflect the dynamics of our system in the first moments and covariances [34]. We can also use Wick’s theorem to contract higher-order moments into these first- and second-order moments. This is useful since we can perform our dynamics with finite-matrix theory without dealing with an exponentially growing Hilbert space. The equation of motion for the operators can then be written in a linearized form as,
where is the Lindbladian and is the dynamical matrix derived from the Heisenberg equations of motion. Assuming a high Q-factor of the cavity, we shall ignore the dissipative dynamics, and consequently the Lindbladian from Eq. (12) in this study. For one pair of signal and idler photons coupled to the cavity and material excitation, the dynamical matrix, , takes the form,
For the second moments, we can define the covariance matrix as,
and the equation of motion for the covariances can be derived to be,
where Eq. (13) has been used. Eq. (17) is known as the Sylvester differential equation and its stationary form is the well-known Lyapunov equation used in the study of chaos [35]. Solving the Sylvester equation allows us to understand the evolution of signal-idler correlations as they interact with a microcavity. In solving this, we must ensure that it concurs with the boundary conditions of the correlation matrix. To establish this, we convert it into its respective input and output forms as follows:
Since the output correlation matrix evolves back in time, we reflect that by switching the hermitian conjugated dynamical matrices in the two terms. Taking the Laplace transform of Eqs. (18) and (19), we get,
where is the correlation matrix of the output. We take so that is the integration of correlations over all time. Since the biphotons are taken to be non-interacting at asymptotic times and interacting at finite times, this time-integrated quantity is effective in capturing the dynamics of the experiment. It evolves under the full Hamiltonian and is useful in setting up our boundary conditions. Fixing , we obtain,
Solving for these equations, we get,
In order to find the output correlation function after all interactions have died out, we need to express in terms of . For this, we go back to Eq. (13) and take its Laplace transform,
where is the dynamical matrix responsible for propagating the vector of moments forward in time from . To connect this to the moments at , we define and propagate the moments backward in time using the Laplace transform,
where we can eliminate the intermediate vector of operators and obtain,
Let us assign . This matrix takes the form of Møller operators as defined in Eqs. (8) - (10). Eq. (27) physically says that the input modes are propagated forward in time and the output modes are propagated backward in time under the interacting Hamiltonian. In doing so, both modes must meet at a common point in time, , to maintain continuity. This forms the essence in which we connect the input and output Heisenberg operators. Since the covariance matrix is defined as,
we can write the transformation between the input and output covariance matrices as,
Now we substitute Eq. (29) in (II.2) to obtain the connection between the input and output covariance matrices as,
Thus, Eq. (30) establishes the connection between the input correlations at with the output correlations at and becomes the governing equation of our method.
In the absence of the signal/idler photon interactions with the cavity, the dynamical matrix represents a free-evolution, , where the off-diagonal signal/idler-cavity coupling . The signal/idler subspace of is then diagonal. Consequently, these photons do not interact with the material degrees of freedom. In this case, the first four terms of Eq. (30) do not contribute to the signal/idler subspace of the output covariance matrix, and thus, the output JSA is identical to the input JSA contained in the initial covariance matrix, . When , the first four terms make a non-diagonal contribution to the subspace of , and we obtain a mapping from under interactions.
This model is also robust to the number of material degrees of freedom present in the cavity as long as the Hamiltonian in Eq. (II.1) generates a gaussian-preserving map. This means that it can be applied equally well to monomeric, dimeric, and other polymeric systems under the two-level approximation. An illustration of such an experimental setup and the theoretical boundary conditions that we described earlier are shown in Figs. 2(a) and 2(b).
II.3 Extension to continuous spectrum
Having laid the theoretical groundwork to describe the experiment, we can now trivially extend this model and create a continuous spectrum of frequency-entangled biphotons. Recognizing that the JSA is a correlation between the signal and idler frequencies, we enter this into the off-diagonal elements of the signal/idler subspace in the extended covariance matrix. The extended dynamical and covariance matrices now look like,
These matrices can now be input in Eq. (30) to obtain the output biphoton correlations.
II.4 Entanglement Entropy
To find the entanglement of the output biphoton states, we perform a Schmidt decomposition of the output JSA.
where are the singular values and are the left and right eigenvalues of the JSA. The von Neumann entropy can then be calculated as,
To derive the purity of the state, we Weyl transform the problem and express the corresponding Wigner function of our multi-mode state in terms of the covariance matrix as descrobed in Ref [34]. This takes the form,
Here, , are the dimensions of the bipartitioned subspaces. We take since local transformations do not change the entanglement structure. Since purity is defined as , we observe that for any operator that admits a well-defined Wigner function, , we can use the overlap property between Wigner functions and write,
This gives us the framework to calculate the purity and, subsequently, the mutual information of the output biphoton state. This Wigner function approach aided by the correlation matrix provides a strong theoretical tool to calculate many other observables as described in Ref [34].
III Results
III.1 Gaussian Initial JSI
To further illustrate our theoretical framework, we begin by considering an input JSI that is a squeezed antisymmetric gaussian state centered at any arbitrary point along the diagonal. The correlation matrix can be trivially written for this state and the resulting output correlation matrix can be obtained using Eq. (30). In Fig. 3, we observe that this does not cause a significant change in the output JSI. This is because a squeezed gaussian corresponds to a state of low entropy. As a result, this leaves little room to bring an observable change in the final JSI from any information transfer into the cavity.
III.2 Experimental Initial JSI
To benchmark the model, we take an experimental input JSI and observe the dynamics. In Fig. 4, we see that the JSI shifts towards higher idler and lower signal wavelengths. Since the system is set up such that the cavity and material dispersions are nearly in resonance with the idler wavelength (), these energies are preferentially absorbed by the cavity and material with a higher probability as compared to the signal photons. As a result, we see a shift in the JSI towards higher idler wavelengths (lower frequencies). For higher coupling strengths, this process occurs at a much faster time scale because of an increase in the Rabi frequency. As a result, for asymptotic output times, the JSI tends to saturate toward higher(lower) idler(signal) wavelengths for higher values of the coupling . The reciprocal behavior between the signal and idler shifts is due to the conservation of energy. We also see a squeezing of the map indicating the filtering effect of cavities. This result is in close agreement with the experimental study conducted in empty microcavities [10] as seen in Fig. 1
Bittner et al. [27] also proved that by measuring the change in the photon entanglement entropy, one can find a direct measure of material correlations. To quantify this, we study its von Neumann entropy given in Eqs. (33) and (34). In Fig. 6, we see that the output entanglement entropy decreases monotonically as a function of the coupling strength, . This occurs because the relatively preferential absorption of the idler photons by the cavity and material modes degrades the correlations between the signal and idler photons. For dimers, the rate of decrease is higher than that for monomers, as seen by the red curve. This is because excitations are able to explore a higher Hilbert space and delocalize into these modes faster in the same amount of time. An intriguing aspect of this plot is that the entropy increases after a certain value of the coupling strength. This points toward a revival of strong correlations in the output signal and idler photons.
Fig. 4 also reveals the emergence of peaks along the off-diagonal, indicating an amplification of signal-idler correlations between specific frequency modes. This behavior can be understood by recognizing that our system comprises two discrete modes—the cavity and material excitations—and two continua of states: the signal and idler modes. To capture the dynamics of a discrete mode coupled to a continuum, one typically discretizes the continua, computes the transition amplitudes, and then takes the continuum limit by letting the discretization spacing approach zero. This procedure enables the calculation of physically relevant quantities, such as level shifts, decay rates, cross sections, and more. A detailed account of this formalism is provided in Cohen-Tannoudji et al. [36] When a discrete mode couples to multiple discrete states that are well isolated from other states in the system, the resulting transition amplitude becomes a superposition of Rabi oscillations with varying frequencies and amplitudes. In the continuum limit, these superpositions lead to an effective irreversible decay of the discrete mode into the continuum. The resulting overlap of the discrete state with the quasi-continuum forms a Lorentzian lineshape centered at the discrete state energy with a width , where denotes the decay rate of the mode.
In our setup, when the discrete cavity mode couples to the signal-idler continua, the cavity excitations decay equally into both continua for the same values of signal-idler frequencies. Because the initial correlations are antisymmetric with respect to the signal and idler, and lie along the diagonal, this decay process leads to an emergence of off-diagonal correlations, manifesting as peaks in Fig. 4. The JSI plotted here is obtained by tracing out the signal-idler submatrix from the final correlation matrix of the complete system and thus reflects the excitations that have decayed into the quasi-continuum.
Additionally, if the material were coupled to the continua in such a way that the ratio of its couplings did not lie in limiting regimes, the resulting output states would exhibit features characteristic of Fano-type asymmetric lineshapes. The trends observed above highlight the capability of photon entanglement to serve as a sensitive probe for exploring many-body interactions and correlations in quantum spectroscopy.
The simplified structure of our model, while not capturing all nonlinear and higher-order effects, enables analytical tractability and provides a lucid method to connect the input and output variables. It successfully reproduces key experimental features and serves as a useful baseline for interpreting the role of entanglement in spectroscopic signals. However, this same formalism necessarily omits certain processes, such as pure dephasing, typically modeled via -type jump operators in the Lindbladian. Stated succinctly, the assumptions of bilinearity and linear coupling prevent the model from capturing non-Markovian dynamics, crosstalk, and higher-order excitations. These effects often arise in experimental systems due to energetic disorder, field inhomogeneity, and cascaded transitions [29]. However, our formalism provides a robust starting point for systematically extending the model to include such interactions. This forms the basis of our future investigation.
IV Summary
Recent advancements in quantum light spectroscopy have revealed exciting possibilities of using entangled photons as sensitive probes of many-body dynamics and correlations in materials. Although experimental and theoretical efforts continue to evolve in parallel, the literature still lacks a framework that accurately explains observed phenomena. The exponential growth of the Hilbert space and the presence of nonlinear and many-body effects make direct simulations computationally prohibitive. To address this, we reformulated the problem using a finite-sized correlation matrix based on a bilinear bosonic Hamiltonian leading to gaussian-preserving time evolutions. This approach allows us to retain the essential physical dynamics that occurring in bipolaritonic systems.
We applied our model to an experimentally measured joint spectral intensity (JSI) of frequency-entangled biphotons generated through spontaneous parametric down conversion. Our simulations predict a shift in the output JSI that is consistent with the experimental observations. Moreover, we observed the emergence of off-diagonal peaks in the JSI, which indicate enhanced correlations between specific signal-idler frequency pairs. We interpret these features as signatures of a discrete cavity mode decaying into the continua of signal and idler modes. This decay redistributes the spectral amplitudes and gives rise to interference effects similar to Fano resonances.
These results underscore the sensitivity of biphoton entanglement to serve as a probe in quantum spectroscopic experiments. Finally, we plan to use this as a theoretical baseline to extend and incorporate more complex interactions in our future work.
V Acknowledgments
The work at the University of Houston was funded by the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337). ARSK acknowledges funding from the National Science Foundation CAREER grant (CHE-2338663), start-up funds from Wake Forest University, funding from the Center for Functional Materials at Wake Forest University.
VI Data Availability Statement
All data and source code used in this work are publicly available in the following GitHub repository: https://github.com/SameerD-phys/Biphoton_entanglement.